ganeshie8 (ganeshie8):

Fun triple integral Evaluate \[\large \iiint_D (xyz)^m(1-x-y-z)^n dxdydz\] \(D : x\ge 0, y\ge 0, z\ge 0, x+y+z\le 1\) \(m,n \in \mathbb{Z^+}\) Answer : \(\dfrac{m!^3 n!}{(3m+n+3)!}\)

OpenStudy (freckles):

@ganeshie8 I never did master setting up triple integrals... but I keep getting 0<x<1-y-z and 0<y<1 since the y intercept of x+y+z=1 is (0,1,0) and 0<z<1 since the z intercept of x+y+z=1 is (0,0,1)

OpenStudy (freckles):

\[\int\limits_{0}^{1} \int\limits_0^1 \int\limits_0^{1-y-z} f(x,y,z) dx dy dz \] but this isn;t working for me

ganeshie8 (ganeshie8):

``` z : 0 -->1 x : 0 --> 1-y-z ``` are looking fine... bounsd for y should be `0-->1-z`

OpenStudy (freckles):

oh ok that gives me the same answer as you have in the end for m=3 and n=2

OpenStudy (freckles):

I was just seeing if I find the same answer for some little m and n

ganeshie8 (ganeshie8):

nice xD i didnt verify the answer for specific cases...

ganeshie8 (ganeshie8):

D represents the solid region below x+y+z=1 plane in first octant

OpenStudy (freckles):

The actual integration is looking nasty.

OpenStudy (freckles):

Don't tell me any hints yet.

ganeshie8 (ganeshie8):

Okay :) I see how it can be nasty... here is a not so useful hint : getting a recurrence relation for the integral \(I(m,n) = \int\limits_0^a x^m(a-x)^n dx\) and evaluating it might reduce the amount of work involved...

OpenStudy (freckles):

\[I=\frac{n}{m+1} \cdot \frac{n-1}{m+2} \cdot \frac{n-2}{m+3} \cdot \frac{n-3}{m+4} \cdots \frac{n-(k-1)}{m+k} \int\limits _0^ax^{m+k} (a-x)^{n-k} dx\]

ganeshie8 (ganeshie8):

Woah! thats fast! i think we need to evaluate the last piece manually and simplify the product..

OpenStudy (freckles):

I needed n steps instead of k

OpenStudy (freckles):

\[I=\frac{n}{m+1} \cdot \frac{n-1}{m+2} \cdot \frac{n-2}{m+3} \cdot \frac{n-3}{m+4} \cdots \frac{n-(k-1)}{m+k} \int\limits\limits _0^ax^{m+k} (a-x)^{n-k} dx \\ I=\frac{n! m!}{(m+n)!}\int\limits_0^a x^{m+n}(a-x)^{0} dx = \frac{n! m!}{(m+n)!} \int\limits_0^a x^{m+n} dx\]

OpenStudy (freckles):

\[I=\frac{n! m!}{(m+n)!} \cdot x^{m+n+1} \frac{1}{m+n+1} |_0^a =\frac{a^{m+n+1}}{(m+n+1)!}n! m!\]

ganeshie8 (ganeshie8):

Nice :) actually thats a beta integral, we worked it without using the beta function xD

OpenStudy (freckles):

I'm glad we didn't that function I don't know it :p

ganeshie8 (ganeshie8):

* \[I(m,n) =\int\limits_0^a x^m(a-x)^n = \dfrac{m!n!}{(m+n+1)!}a^{m+n+1}\]

OpenStudy (freckles):

\[ \int\limits_{0}^{1-y-z}(xym)^m(1-y-z=z)^n dx \\ (ym)^m \int\limits_0^{1-y-z}x^m(1-y-z-x)^n dx \\ \text{ let } a=1-y-z \\ =(ym)^m \cdot \frac{m! n!}{(m+n+1)!} (1-y-z)^{m+n+1} \]

OpenStudy (freckles):

now I have to do the dy and then dz thingy

OpenStudy (freckles):

it the dy part should be very similar

ganeshie8 (ganeshie8):

looks neat, is there a typo... \((xy\color{Red}{z})^m\)

OpenStudy (freckles):

yes that particular m was suppose to be z

OpenStudy (freckles):

\[ \int\limits_{0}^{1-y-z}(xyz)^m(1-y-z-x)^n dx \\ (yz)^m \int\limits_0^{1-y-z}x^m(1-y-z-x)^n dx \\ \text{ let } a=1-y-z \\ =(yz)^m \cdot \frac{m! n!}{(m+n+1)!} (1-y-z)^{m+n+1} \] corrections made even that equal sign thingy

OpenStudy (freckles):

and that other z thingy lol

OpenStudy (freckles):

\[\frac{z^m m! n!}{(m+n+1)!} \int\limits_0^{1-z} y^m (1-z-y)^{m+n+1} dy\]

OpenStudy (freckles):

I know our a here is going to be the 1-z but the exponent thing is going to be a little bit different I think I just need to replace our old n with m+n+1

OpenStudy (freckles):

\[ \frac{z^m m! n!}{(m+n+1)!} \frac{m! (m+n+1)!}{(m+(m+n+1)+1)!} (1-z)^{m+(m+n+1)+1}\]

OpenStudy (freckles):

and I still need to do the last integral part

OpenStudy (freckles):

let me know if i made mistake in saying what I said

ganeshie8 (ganeshie8):

\[\frac{z^m m! n!}{(m+n+1)!} \frac{m! (m+n+1)!}{(m+(m+n+1)+1)!} (1-z)^{m+(m+n+1)+1} \\~\\~\\= \dfrac{z^mm!^2 n!}{(2m+n+2)!} (1-z)^{2m+n+2}\] looks good to me..

OpenStudy (freckles):

ok just checking because I feel like I overused that brain thing

OpenStudy (freckles):

\[\frac{ m! n!}{(m+n+1)!} \frac{m!(m+n+1)!}{(2m+n+2)!} \int\limits\limits_0^1 z^m (1-z)^{2m+n+2} dz \\ \]

OpenStudy (freckles):

\[\frac{m! n!}{(m+n+1)!} \frac{m! (m+n+1)!}{(2m+n+2)!} \frac{m! (2m+n+2)!}{(m+(2m+n+2)+1)!} =\frac{m!^3n!}{(3m+n+3)!}\] wowzers

OpenStudy (freckles):

I don't think I would have been able to do it without your "useless hint"

ganeshie8 (ganeshie8):

Sweeet :) nothing feels better than seeing a hard problem solved neatly xD

OpenStudy (freckles):

Nothing feels greater than feeling like you solved a problem you couldn't solve before thanks for the problem

OpenStudy (freckles):

as in me like I don't think I would have attempted to evaluate the problem and if i did try I think I would have given up

ganeshie8 (ganeshie8):

From your solution it seems below is true \[\large \iiint_D x^ay^bz^c(1-x-y-z)^n dxdydz = \dfrac{a!b!c!n!}{(a+b+c+n+3)!}\] not sure... wil try some other time :)

OpenStudy (freckles):

yeah I think I will have to look at that later too because I don't think I can use my brain anymore

OpenStudy (freckles):

and it does kinda look it that is true

OpenStudy (freckles):

I will do one small test and see

OpenStudy (dan815):

im not sure but i think from the integral i can tell theres a separable form to it?

OpenStudy (dan815):

which wed get out of the binomialexpansion

OpenStudy (freckles):

it does work for a=2,b=1c=3,n=4 and I know this doesn't prove for all a,b,c,n in the positive integers

OpenStudy (dan815):